Steam at T = 300C flows in a cast iron pipe shell with thermal conductivity k1=80W/mC, inner and outer diameter D1 = 5cm, and D2 =6cm respectively, and length L = 1m (notice that the cast iron pipe shell is a zone of conduction physically realized by the space between two concentric cylinders). The heat transfer coefficient for the steam inside the pipe is h1 = 60 W/m2. Furthermore, heat is lost to the surroundings at T2 = 5C through convection with a heat transfer coefficient h2 = 20 W/m2. Determine the rate of heat loss and the temperature drop across the pipe shell (assume steady state conditions and absence of chemical/nuclear reactions).
Answer #1

kz Bowlmk DI= 5cm do 2.5cm. D₂ = bein, d. = 8cm L=1m 300°c sowik zow/w 7 =500 +273 2573K T = 5+ 273-278K m - T.D, L = T (0.05